OpenStudy (cometailcane):

Solve for x. sin^-1(1-3x)=-pi/6

OpenStudy (idku):

\(\color{black}{\displaystyle \sin^{-1}(1-3x)=-\pi/6 }\) \(\color{black}{\displaystyle \sin^{}(-\pi/6)=1-3x }\) \(\color{black}{\displaystyle -\sin^{}(\pi/6)=1-3x }\) \(\color{black}{\displaystyle -\sin^{}(30^\circ)=1-3x }\) \(\color{black}{\displaystyle -\sin^{}(30^\circ)-1=-3x }\) \(\color{black}{\displaystyle (-\sin^{}(30^\circ)-1)\div (-3)=x }\)

OpenStudy (idku):

just follow the basic steps ...